OpenStudy (anonymous):

if r and s are constant and x^2+rx+12 is equivalent to (x+3)(x+s) what is the value of r?

OpenStudy (zzr0ck3r):

(x+3)(x+s) = x^2+x(s+3)+3s = x^2+rx+12 3s=12 implies s=4 so 7x=rx implies r = 7